# Zero-Shot Neural Architecture Search with Weighted Response Correlation

### v1

@@ 4.2 Analysis of Proxy Evaluation
Table 4: The comparison of different proxies across the CIFAR-10 (c10), CIFAR-100 (c100), and ImageNet16-120 (im120) datasets.
Proxies PNorm PGNorm SNIP GraSP SynFlow Fisher JacCor
c10 0.684 0.591 0.593 0.543 0.743 0.506 0.726
c100 0.712 0.648 0.648 0.555 0.771 0.563 0.721
im120 0.711 0.588 0.588 0.551 0.755 0.501 0.716
avg. 0.702 0.609 0.610 0.550 0.756 0.523 0.721
std. 0.013 0.028 0.027 0.005 0.011 0.028 0.004
Proxies EPE Phi Zen BT BndlEnt LyDynIso DynIso
c10 0.693 0.703 0.257 0.510 0.199 0.633 0.662
c100 0.602 0.730 0.297 0.449 0.192 0.682 0.728
im120 0.566 0.663 0.293 0.460 0.130 0.668 0.674
avg. 0.620 0.699 0.282 0.473 0.174 0.661 0.688
std. 0.053 0.028 0.018 0.027 0.031 0.021 0.029
Proxies NetSens NASWOT ZiCo Params FLOPs
c10 0.686 0.779 0.800 0.750 0.730
c100 0.706 0.818 0.810 0.730 0.710
im120 0.624 0.791 0.790 0.690 0.670
avg. 0.672 0.796 0.800 0.723 0.703
std. 0.035 0.016 0.008 0.025 0.025
Proxies ACor RCor WRCor ZeroCost SPW SJW ValAcc
c10 0.787 0.803 0.811 0.775 0.789 0.851 0.842
c100 0.823 0.839 0.846 0.802 0.823 0.846 0.845
im120 0.790 0.817 0.822 0.775 0.808 0.847 0.833
avg. 0.800 0.820 0.826 0.784 0.807 0.848 0.840
std. 0.016 0.015 0.015 0.013 0.014 0.002 0.005
@@
Table 5: The comparison of different proxies across different batch sizes of 32, 64, and 128.
Proxies PNorm PGNorm SNIP GraSP SynFlow Fisher JacCor
32 0.684 0.597 0.599 0.522 0.743 0.510 0.730
64 0.684 0.591 0.593 0.543 0.743 0.506 0.726
128 0.684 0.586 0.589 0.519 0.743 0.507 0.735
avg. 0.684 0.591 0.594 0.528 0.743 0.508 0.730
std. 0.000 0.004 0.004 0.011 0.000 0.002 0.004
Proxies EPE Phi Zen BT BndlEnt LyDynIso DynIso
32 0.661 0.701 0.258 0.466 0.185 0.633 0.654
64 0.693 0.703 0.257 0.510 0.199 0.633 0.662
128 0.703 0.704 0.256 0.541 0.228 0.633 0.667
avg. 0.686 0.703 0.257 0.506 0.204 0.633 0.661
std. 0.018 0.001 0.001 0.031 0.018 0.000 0.005
Proxies NetSens NASWOT ZiCo Params FLOPs
32 0.686 0.775 - 0.750 0.730
64 0.686 0.779 - 0.750 0.730
128 0.686 0.782 - 0.750 0.730
avg. 0.686 0.779 - 0.750 0.730
std. 0.000 0.003 - 0.000 0.000
Proxies ACor RCor WRCor ZeroCost SPW SJW ValAcc
32 0.781 0.791 0.797 0.783 0.786 0.829 0.842
64 0.787 0.803 0.811 0.775 0.789 0.851 0.842
128 0.780 0.812 0.818 0.768 0.791 0.814 0.842
avg. 0.783 0.802 0.809 0.775 0.789 0.831 0.842
std. 0.003 0.009 0.009 0.006 0.002 0.015 0.000
@@
Table 6: The comparison of different proxies across Kaiming uniform (uniform), Kaiming normal (normal), and $N(0,1)$ (n01) initializations.
Proxies PNorm PGNorm SNIP GraSP SynFlow Fisher JacCor
uniform 0.684 0.591 0.593 0.543 0.743 0.506 0.726
normal 0.692 0.559 0.570 0.451 0.738 0.480 0.739
n01 0.725 0.369 0.518 0.106 0.743 0.453 0.733
avg. 0.700 0.506 0.560 0.367 0.741 0.480 0.733
std. 0.018 0.098 0.031 0.188 0.002 0.022 0.005
Proxies EPE Phi Zen BT BndlEnt LyDynIso DynIso
uniform 0.693 0.703 0.257 0.510 0.199 0.633 0.662
normal 0.698 0.670 0.036 0.558 0.028 0.612 0.587
n01 0.719 0.657 0.352 0.568 0.040 0.717 0.358
avg. 0.703 0.677 0.215 0.545 0.089 0.654 0.536
std. 0.011 0.019 0.132 0.025 0.078 0.045 0.129
Proxies NetSens NASWOT ZiCo Params FLOPs
uniform 0.686 0.779 - 0.750 0.730
normal 0.641 0.777 0.800 0.750 0.730
n01 0.625 0.777 - 0.750 0.730
avg. 0.651 0.778 - 0.750 0.730
std. 0.026 0.001 - 0.000 0.000
Proxies ACor RCor WRCor ZeroCost SPW SJW ValAcc
uniform 0.787 0.803 0.811 0.775 0.789 0.851 0.842
normal 0.775 0.798 0.805 0.791 0.780 0.829 0.842
n01 0.766 0.794 0.796 0.700 0.804 0.839 0.842
avg. 0.776 0.798 0.804 0.755 0.791 0.840 0.842
std. 0.009 0.004 0.006 0.040 0.010 0.009 0.000
@@
Table 7: The comparison of different proxies across NAS-Bench-101 (NB101) and NAS-Bench-201 (NB201) search spaces. $|\Delta|$: the absolute deviation of the Spearman $\rho$ values.
Proxies PNorm PGNorm SNIP GraSP SynFlow Fisher JacCor
NB101 0.561 0.248 0.157 0.357 0.429 0.281 0.362
NB201 0.684 0.591 0.593 0.543 0.743 0.506 0.726
$|\Delta|$ 0.123 0.343 0.436 0.186 0.314 0.225 0.364
Proxies EPE Phi Zen BT BndlEnt LyDynIso DynIso
NB101 0.008 0.016 0.200 0.054 0.012 0.486 0.281
NB201 0.693 0.703 0.257 0.510 0.199 0.633 0.662
$|\Delta|$ 0.685 0.687 0.057 0.456 0.187 0.147 0.381
Proxies NetSens NASWOT ZiCo Params FLOPs
NB101 0.302 0.410 - 0.440 0.430
NB201 0.686 0.779 0.800 0.750 0.730
$|\Delta|$ 0.384 0.369 - 0.310 0.300
Proxies ACor RCor WRCor ZeroCost SPW SJW ValAcc
NB101 0.536 0.553 0.534 0.000 0.569 0.383 0.647
NB201 0.787 0.803 0.811 0.775 0.789 0.851 0.842
$|\Delta|$ 0.251 0.250 0.277 0.775 0.220 0.468 0.195
How Important Is the Input for Our Proxies?
@@
How Do They Perform for Top-Ranked Architectures?
In the NAS task, we usually focus more on architectures with higher validation accuracy, as the goal of NAS is to discover the optimal architecture. Table 8 reports the Spearman $\rho$ values of all mentioned proxies calculated using the top-20% superior architectures from the NAS-Bench-201 search space. The Spearman $\rho$ values of all proxies significantly decrease when considering these top-ranked architectures. Although WRCor performs worse than SynFlow and BT, it remains competitive. Therefore, for the top-ranked architectures, our proposed proxies have some competitiveness but a significant reduction in terms of Spearman $\rho$ value.
Table 8: The Spearman $\rho$ values of different proxies for the top-20% superior architectures. $|\Delta|$: the absolute deviation of the Spearman $\rho$ values.
Proxies PNorm PGNorm SNIP GraSP SynFlow Fisher JacCor
1000 0.684 0.591 0.593 0.543 0.743 0.506 0.726
top20% 0.146 0.110 0.105 0.078 0.466 0.133 0.126
$|\Delta|$ 0.538 0.481 0.488 0.465 0.277 0.373 0.600
Proxies EPE Phi Zen BT BndlEnt LyDynIso DynIso
1000 0.693 0.703 0.257 0.510 0.199 0.633 0.662
top20% 0.073 0.044 0.134 0.242 0.190 0.094 0.026
$|\Delta|$ 0.620 0.659 0.123 0.268 0.009 0.539 0.636
Proxies NetSens NASWOT ZiCo Params FLOPs
1000 0.686 0.779 0.800 0.750 0.730
top20% 0.073 0.124 - - -
$|\Delta|$ 0.613 0.655 - - -
Proxies ACor RCor WRCor ZeroCost SPW SJW ValAcc
1000 0.787 0.803 0.811 0.775 0.789 0.851 0.842
top20% 0.113 0.122 0.158 0.024 0.282 0.462 0.440
$|\Delta|$ 0.674 0.681 0.653 0.751 0.507 0.389 0.402
@@
Some NAS works [16, 30] demonstrate that we can achieve an omnipotent architecture performance predictor by integrating multiple proxies and predictors. To capitalize on the strengths of different proxies to improve their Spearman $\rho$ values for the above challenging scenarios, such as evaluating the top-ranked architectures and exploring the large search space, we employ a majority voting rule among various proxies for architecture estimation. Although all mentioned voting proxies rely on SynFlow, there are many differences between them, as shown in Table 2, 4, 5, 6, 7, and 8. SPW and SJW that utilize WRCor usually have significantly higher Spearman $\rho$ value, which demonstrates that WRCor is a crucial and effective proxy. Since JacCor performs better than PNorm, SJW outperforms SPW in most scenarios, except when exploring the larger NAS-Bench-101 search space. The reason for the performance reduction of SJW in the larger NAS-Bench-101 search space is attributed to the fact that JacCor [13] estimates networks using the histograms ($0<(\Sigma_{J})_{i,j}<\beta$) of the correlation coefficient matrices. This is a coarse-grained score and can lead to similar architectures receiving the same scores as the search space expands unless the upper bound $\beta$ is appropriately adjusted. SPW is a better proxy when searching in the larger search space. And SJW has the highest Spearman $\rho$ value on top-ranked architectures. In summary, SPW and SJW compensate for the shortcomings of our proposed proxies in two challenging scenarios, which can ensure a more accurate architecture evaluation.
@@ 4.3.1 Search on NAS-Bench-101 Search Space
To compare our zero-shot NAS algorithms across various proxies and search strategies, we conduct experiments on the NAS-Bench-101 benchmark. Table 9 reports their validation/test accuracies and search costs on the CIFAR-10 dataset. Compared with X-ZeroCost, X-WRCor promotes the test accuracies by $1.99\sim 8.81$ with higher robustness. RL-WRCor and R-WRCor discover the second and third competitive architectures with the test accuracies of 93.36% and 93.26%, respectively. X-SJW slightly outperforms X-WRCor. Since X-SJW replaces SNIP with WRCor for voting, it has the $2.10\sim 8.86$ higher test accuracies with higher robustness than X-ZeroCost. Notably, RL-SJW discovers the most competitive architectures with an average test accuracy of 93.47% and a minimal standard deviation of 0.03. Furthermore, X-WRCor and X-SJW discover architectures that are comparable to those discovered by X-ValACC, but X-ValACC requires a $25\sim 50$ higher cost. Our proposed WRCor and SJW proxies are effective, efficient, and robust estimation strategies for NAS.
Table 9: The comparison of zero-shot NAS across different search strategies and training-free proxies on the NAS-Bench-101 benchmark. We run each algorithm three times where $N=1000$. , , and represent the top-3 algorithms in the last three blocks. X-Y represents the zero-shot NAS algorithm that uses search strategy X and training-free proxy Y.
@@ 4.3.2 Search on NAS-Bench-201 Search Space
To further compare the existing zero-shot NAS algorithms with ours, we conduct experiments on the NAS-Bench-201 benchmark. Table 10 reports their test accuracies and search costs on the CIFAR-10, CIFAR-100, and ImageNet16-120 datasets. X-WRCor outperforms most existing NAS algorithms, except GDAS, TE-NAS, and ZiCo on CIFAR-10 and CIFAR-100 datasets. RE-SJW significantly outperforms most algorithms, except TE-NAS and ZiCo on the CIFAR-10 dataset and RE-ValACC. This discrepancy can be because our proxies offer a more accurate reflection of real-world performance, especially on datasets with numerous categories. Another reason is that the datasets with fewer categories tend to have samples within a minibatch that are more similar, which can potentially disturb our proxies and affect their performance. In essence, our proposed proxies are more suitable for datasets with a large number of categories. Besides, RE-WRCor has a significant performance reduction compared to R-WRCor and RL-WRCor across the three datasets. This is probably because our algorithm with WRCor can easily obtain inferior local optima. SJW can leverage the strengths of various proxies to improve the solution quality. RE-SJW is a promising zero-shot NAS algorithm in terms of performance and efficiency, achieving competitive architectures with up to 50 times efficiency compared to X-ValAcc and existing NAS algorithms.
Table 10: The comparison of zero-shot NAS across different search strategies and training-free proxies on the NAS-Bench-201 benchmark. We run each algorithm three times where $N=1000$.
@@ 4.3.3 Search on MobileNetV2 Search Space
We also conduct experiments of architecture search on the ImageNet-1k dataset in the MobileNetV2 search space. We employ the optimal RE-SJW according to the experimental results on the NAS-Bench-101/201 benchmarks. Table 11 reports the test error on the ImageNet-1k dataset, the number of parameters, and the number of FLOPs for our discovered architecture below 600M FLOPs, as well as the search cost. The discovered architecture by RE-SJW is detailed in Table 12. This architecture achieves a competitive result of 22.1% top-1 test error on the ImageNet-1k dataset in four GPU hours, thus being more efficient than existing NAS algorithms. It outperforms all manually designed architectures and most architectures designed by existing NAS algorithms. Compared with existing zero-shot NAS algorithms, our discovered architecture achieves 2.4% less test error than TE-NAS with the same search time. We discover a comparable architecture with significantly less search cost. Therefore, RE-SJW is an effective and efficient zero-shot NAS algorithm.
Table 11: The comparison to the discovered architectures below 600M FLOPs by different NAS algorithms on the ImageNet-1k dataset.
Test Error (%) Params FLOPs Cost

### v2

@@ 4.2 Analysis of Proxy Evaluation
Table 4: The comparison of different proxies across the CIFAR-10 (c10), CIFAR-100 (c100), and ImageNet16-120 (im120) datasets on the NAS-Bench-201 benchmark.
Proxies c10 c100 im120 avg. std.
PNorm 0.684 0.712 0.711 0.702 0.013
PGNorm 0.591 0.648 0.588 0.609 0.028
SNIP 0.593 0.648 0.588 0.610 0.027
GraSP 0.543 0.555 0.551 0.550 0.005
SynFlow 0.743 0.771 0.755 0.756 0.011
Fisher 0.506 0.563 0.501 0.523 0.028
JacCor 0.726 0.721 0.716 0.721 0.004
EPE 0.693 0.602 0.566 0.620 0.053
Phi 0.703 0.730 0.663 0.699 0.028
Zen 0.257 0.297 0.293 0.282 0.018
BT 0.510 0.449 0.460 0.473 0.027
BndlEnt 0.199 0.192 0.130 0.174 0.031
LyDynIso 0.633 0.682 0.668 0.661 0.021
DynIso 0.662 0.728 0.674 0.688 0.029
NetSens 0.686 0.706 0.624 0.672 0.035
NASWOT 0.779 0.818 0.791 0.796 0.016
ZiCo 0.800 0.810 0.790 0.800 0.008
Params 0.750 0.730 0.690 0.723 0.025
FLOPs 0.730 0.710 0.670 0.703 0.025
ValAcc 0.842 0.845 0.833 0.840 0.005
ACor (Ours) 0.787 0.823 0.790 0.800 0.016
RCor (Ours) 0.803 0.839 0.817 0.820 0.015
WRCor (Ours) 0.811 0.846 0.822 0.826 0.015
ZeroCost 0.775 0.802 0.775 0.784 0.013
SPW (Ours) 0.789 0.823 0.808 0.807 0.014
SJW (Ours) 0.851 0.846 0.847 0.848 0.002
@@
Table 5: The comparison of different proxies across different batch sizes of 32, 64, and 128 on the CIFAR-10 dataset and the NAS-Bench-201 benchmark.
Proxies 32 64 128 avg. std.
PNorm 0.684 0.684 0.684 0.684 0.000
PGNorm 0.597 0.591 0.586 0.591 0.004
SNIP 0.599 0.593 0.589 0.594 0.004
GraSP 0.522 0.543 0.519 0.528 0.011
SynFlow 0.743 0.743 0.743 0.743 0.000
Fisher 0.510 0.506 0.507 0.508 0.002
JacCor 0.730 0.726 0.735 0.730 0.004
EPE 0.661 0.693 0.703 0.686 0.018
Phi 0.701 0.703 0.704 0.703 0.001
Zen 0.258 0.257 0.256 0.257 0.001
BT 0.466 0.510 0.541 0.506 0.031
BndlEnt 0.185 0.199 0.228 0.204 0.018
LyDynIso 0.633 0.633 0.633 0.633 0.000
DynIso 0.654 0.662 0.667 0.661 0.005
NetSens 0.686 0.686 0.686 0.686 0.000
NASWOT 0.775 0.779 0.782 0.779 0.003
ZiCo - - - - -
Params 0.750 0.750 0.750 0.750 0.000
FLOPs 0.730 0.730 0.730 0.730 0.000
ValAcc 0.842 0.842 0.842 0.842 0.000
ACor (Ours) 0.781 0.787 0.780 0.783 0.003
RCor (Ours) 0.791 0.803 0.812 0.802 0.009
WRCor (Ours) 0.797 0.811 0.818 0.809 0.009
ZeroCost 0.783 0.775 0.768 0.775 0.006
SPW (Ours) 0.786 0.789 0.791 0.789 0.002
SJW (Ours) 0.829 0.851 0.814 0.831 0.015
@@
Table 6: The comparison of different proxies across Kaiming uniform (uniform), Kaiming normal (normal), and $N(0,1)$ (n01) initializations on the CIFAR-10 dataset and the NAS-Bench-201 benchmark.
Proxies uniform normal n01 avg. std.
PNorm 0.684 0.692 0.725 0.700 0.018
PGNorm 0.591 0.559 0.369 0.506 0.098
SNIP 0.593 0.570 0.518 0.560 0.031
GraSP 0.543 0.451 0.106 0.367 0.188
SynFlow 0.743 0.738 0.743 0.741 0.002
Fisher 0.506 0.480 0.453 0.480 0.022
JacCor 0.726 0.739 0.733 0.733 0.005
EPE 0.693 0.698 0.719 0.703 0.011
Phi 0.703 0.670 0.657 0.677 0.019
Zen 0.257 0.036 0.352 0.215 0.132
BT 0.510 0.558 0.568 0.545 0.025
BndlEnt 0.199 0.028 0.040 0.089 0.078
LyDynIso 0.633 0.612 0.717 0.654 0.045
DynIso 0.662 0.587 0.358 0.536 0.129
NetSens 0.686 0.641 0.625 0.651 0.026
NASWOT 0.779 0.777 0.777 0.778 0.001
ZiCo - 0.800 - - -
Params 0.750 0.750 0.750 0.750 0.000
FLOPs 0.730 0.730 0.730 0.730 0.000
ValAcc 0.842 0.842 0.842 0.842 0.000
ACor (Ours) 0.787 0.775 0.766 0.776 0.009
RCor (Ours) 0.803 0.798 0.794 0.798 0.004
WRCor (Ours) 0.811 0.805 0.796 0.804 0.006
ZeroCost 0.775 0.791 0.700 0.755 0.040
SPW (Ours) 0.789 0.780 0.804 0.791 0.010
SJW (Ours) 0.851 0.829 0.839 0.840 0.009
@@
Table 7: The comparison of different proxies across NAS-Bench-101 (NB101) and NAS-Bench-201 (NB201) search spaces on the CIFAR-10 dataset. $|\Delta|$: the absolute deviation of the Spearman $\rho$ values.
Proxies NB101 NB201 $|\Delta|$
PNorm 0.561 0.684 0.123
PGNorm 0.248 0.591 0.343
SNIP 0.157 0.593 0.436
GraSP 0.357 0.543 0.186
SynFlow 0.429 0.743 0.314
Fisher 0.281 0.506 0.225
JacCor 0.362 0.726 0.364
EPE 0.008 0.693 0.685
Phi 0.016 0.703 0.687
Zen 0.200 0.257 0.057
BT 0.054 0.510 0.456
BndlEnt 0.012 0.199 0.187
LyDynIso 0.486 0.633 0.147
DynIso 0.281 0.662 0.381
NetSens 0.302 0.686 0.384
NASWOT 0.410 0.779 0.369
ZiCo - 0.800 -
Params 0.440 0.750 0.310
FLOPs 0.430 0.730 0.300
ValAcc 0.647 0.842 0.195
ACor (Ours) 0.536 0.787 0.251
RCor (Ours) 0.553 0.803 0.250
WRCor (Ours) 0.534 0.811 0.277
ZeroCost 0.000 0.775 0.775
SPW (Ours) 0.569 0.789 0.220
SJW (Ours) 0.383 0.851 0.468
Does the input of our proposed proxy require relevance to the target dataset?
@@
Some NAS works [19, 33] demonstrate that we can achieve an omnipotent architecture performance predictor by integrating multiple proxies and predictors. To capitalize on the strengths of different proxies to improve their Spearman $\rho$ values for the above challenging scenarios, such as exploring the large search space, we employ a majority voting rule among various proxies for architecture estimation. Although all mentioned voting proxies rely on SynFlow, there are many differences between them, as shown in Table 2, 4, 5, 6, and 7. SPW and SJW that utilize WRCor usually have significantly higher Spearman $\rho$ value, which demonstrates that WRCor is a crucial and effective proxy. Since JacCor performs better than PNorm, SJW outperforms SPW in most scenarios, except when exploring the larger NAS-Bench-101 search space. The reason for the performance reduction of SJW in the larger NAS-Bench-101 search space is attributed to the fact that JacCor [16] estimates networks using the histograms ($0<(\Sigma_{J})_{i,j}<\beta$) of the correlation coefficient matrices. This is a coarse-grained score and can lead to similar architectures receiving the same scores as the search space expands unless the upper bound $\beta$ is appropriately adjusted. SPW is a better proxy when searching in the larger search space. In summary, SPW and SJW compensate for the shortcomings of our proposed proxies , which can ensure a more accurate architecture evaluation.
@@ 4.3.1 Search on NAS-Bench-101 Search Space
To compare our zero-shot NAS algorithms across various proxies and search strategies, we conduct experiments on the NAS-Bench-101 benchmark. Table 8 reports their validation/test accuracies and search costs on the CIFAR-10 dataset. Compared with X-ZeroCost, X-WRCor promotes the test accuracies by $1.99\sim 8.81$ with higher robustness. RL-WRCor and R-WRCor discover the second and third competitive architectures with the test accuracies of 93.36% and 93.26%, respectively. X-SJW slightly outperforms X-WRCor. Since X-SJW replaces SNIP [24] with WRCor for voting, it has the $2.10\sim 8.86$ higher test accuracies with higher robustness than X-ZeroCost. Notably, RL-SJW discovers the most competitive architectures with an average test accuracy of 93.47% and a minimal standard deviation of 0.03. Furthermore, X-WRCor and X-SJW discover architectures that are comparable to those discovered by X-ValACC, but X-ValACC requires a $25\sim 50$ higher cost. Our proposed WRCor and SJW proxies are effective, efficient, and robust estimation strategies for NAS.
Table 8: The comparison of zero-shot NAS across different search strategies and training-free proxies on the CIFAR-10 dataset and the NAS-Bench-101 benchmark. We run each algorithm three times where $N=1000$. , , and represent the top-3 algorithms in the last three blocks. X-Y represents the zero-shot NAS algorithm that uses search strategy X and training-free proxy Y.
@@ 4.3.2 Search on NAS-Bench-201 Search Space
To further compare the existing zero-shot NAS algorithms with ours, we conduct experiments on the NAS-Bench-201 benchmark. Table 9 reports their test accuracies and search costs on the CIFAR-10, CIFAR-100, and ImageNet16-120 datasets. X-WRCor outperforms most existing NAS algorithms, except GDAS, TE-NAS, and ZiCo on CIFAR-10 and CIFAR-100 datasets. RE-SJW significantly outperforms most algorithms, except TE-NAS [23] and ZiCo [21] on the CIFAR-10 dataset, AZ-NAS [41] on the ImageNet16-120 dataset, and RE-ValACC. RE-SJW exhibits superior performance on the CIFAR-100 and ImageNet16-120 datasets compared to the CIFAR-10 dataset. This can be because our proxies offer a more accurate reflection of real-world performance, especially on datasets with numerous categories. Another reason is that the datasets with fewer categories tend to have samples within a minibatch that are more similar, which can potentially disturb our proxies and affect their performance. In essence, our proposed proxies are more suitable for datasets with a large number of categories. Besides, RE-WRCor has a significant performance reduction compared to R-WRCor and RL-WRCor across the three datasets. This is probably because our algorithm with WRCor can easily obtain inferior local optima. SJW can leverage the strengths of various proxies to improve the solution quality. RE-SJW is a promising zero-shot NAS algorithm in terms of performance and efficiency, achieving competitive architectures with up to 50 times efficiency compared to X-ValAcc and existing NAS algorithms.
Table 9: The comparison of zero-shot NAS across different search strategies and training-free proxies on the NAS-Bench-201 benchmark, including CIFAR-10, CIFAR-100, and ImageNet16-120 datasets. We run each algorithm three times where $N=1000$. , , and represent the top-3 algorithms in the last three blocks. X-Y represents the zero-shot NAS algorithm that uses search strategy X and training-free proxy Y.
@@ 4.3.3 Search on MobileNetV2 Search Space
We also conduct experiments of architecture search on the ImageNet-1k dataset in the MobileNetV2 search space. We employ the optimal RE-SJW according to the experimental results on the NAS-Bench-101/201 benchmarks. Table 10 reports the test error on the ImageNet-1k dataset, the number of parameters, and the number of FLOPs for our discovered architecture below 600M FLOPs, as well as the search cost. The discovered architecture by RE-SJW is detailed in Table 11. This architecture achieves a competitive result of 22.1% top-1 test error on the ImageNet-1k dataset in four GPU hours, thus being more efficient than existing NAS algorithms. It outperforms all manually designed architectures and most architectures designed by existing NAS algorithms. Compared with existing zero-shot NAS algorithms, our discovered architecture achieves 2.4% less test error than TE-NAS [23] with the same search time and only 0.2% and 0.5% more test error than ZiCo [21] and AZ-NAS [41], respectively, while requiring less search time. We discover a comparable architecture with significantly less search cost. Therefore, RE-SJW is an effective and efficient zero-shot NAS algorithm.
Table 10: The comparison to the discovered architectures below 600M FLOPs by different NAS algorithms on the ImageNet-1k dataset. Algo. Type refers to the type of the NAS algorithm, including manually designing (M), multi-shot NAS (MS), one-shot NAS (OS), zero-shot NAS (ZS).
Test Error (%) Params FLOPs Cost Algo.